quantum simple harmonic oscillator consisting of an electron under the influence of an restoring force was found to make a transition from n=2 state to n=3 state after absorbing a photon of wavelength 550 nm. Find the proportionality constant k for the oscillator. Mass of an electron =9.1x10-31 Kg; Planck’s constant=6.63x10-34 Js; speed of light in vacuum=3.0x108 m/s
1
Expert's answer
2020-05-24T18:02:10-0400

"E=k(n+0.5)"

"E_2-E_1=1.5k=\frac{hc}{\lambda}"

"1.5k=\frac{(6.63\cdot10^{-34})(3\cdot10^{8})}{550\cdot10^{-9}}"

"k=2.41\cdot10^{-19}J"